OpenStudy (anonymous):

Medal for help :) 6.5 = -3 + n A. n=3.5 B. n=9.5 C. n=19.5 D. n=-19.5

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Just add 3 to both sides 6.5+3=9.5 n=9.5

OpenStudy (jgirl128):

6.5 = -3 + n +3 +3 ___ ____ 9.5 = n n = 9.5

jimthompson5910 (jim_thompson5910):

Plug n = -19.5 into the original equation and show us what you get

OpenStudy (jgirl128):

@JuniorEinstein1 try the equation again, by yourself and see what you get. if you get -19.5 again, do what jim said and plug -19.5 into the original equation.